Show a complete solution to each problem.
Find the exact length of the side of a square whose diagonal is 3 feet.
Draw a square. Draw a diagonal. Let each side length be x.
I don't understand what they are asking and am not sure how to solve this problem. Can someone help explain. Thanks.

Found 2 solutions by rfer, unlockmath:
Answer by rfer(16322)

your using a^2+b^2=c^2
a^2+b^2=9
4.5+4.5=9
sq rt of 4.5=2.12 ft each side
you put it in the form you need for class

Answer by unlockmath(1688)

Hello,
Let's see if I can help you. Basically, we want to find the distance of the sides. Since it's a square both sides are the same. We can use the pythagornean formula to solve this. It's stated this way a^2+b^2=c^2
c is the diagonal and a,b are the legs of a triangle. We'll replace a,b with x since both sides are same length.
x^2+x^2=3^2 Combine like terms gives us:
2x^2=9 Divide both sides by 2
x^2=9/2 Now square root both side so we have x all alone.
x=sq rt of 9/2 This give us
x=3/sq rt of 2
Sq rt of 2 is approx =1.41 then we can do 3 divided by 1.41 = 2.13 (rounded off)
There we go the sides of the sq are 2.13 feet.
RJ Toftness
www.math-unlock.com